OpenStudy (anonymous):

Solve the equation for x. –x + 10 – 8x = 17 – 2x

OpenStudy (marissalovescats):

And and subtract like terms until your x is on the left and the number is on the right then solve.

OpenStudy (souvik):

place the terms with x in one side first

OpenStudy (anonymous):

–x + 10 – 8x = 17 – 2x i.e. 2x -9x = 17-10 i.e. -7x = 7 i.e. x = 7/-7=-1

OpenStudy (marissalovescats):

Yup ^